### v1

# ContraGen: A Multi-Agent Generation Framework for Enterprise Contradictions Detection

Abstract
Retrieval-Augmented Generation (RAG) integrates LLMs with external sources, offering advanced capabilities for information access and decision-making. However, contradictions in retrieved evidence can result in inconsistent or untrustworthy outputs, which is especially problematic in enterprise settings where compliance, governance, and accountability are critical. Existing benchmarks for contradiction detection are limited to sentence-level analysis and do not capture the complexity of enterprise documents such as contracts, financial filings, compliance reports, or policy manuals. To address this limitation, we propose ContraGen, a contradiction-aware benchmark framework tailored to enterprise domain. The framework generates synthetic enterprise-style documents with embedded contradictions, enabling systematic evaluation of both intra-document and cross-document consistency. Automated contradiction mining is combined with human-in-the-loop validation to ensure high accuracy. Our contributions include generating realistic enterprise documents, modeling a taxonomy of contradiction types common in business processes, enabling controlled creation of self- and pairwise contradictions, developing a contradiction-aware retrieval evaluation pipeline and embedding human oversight to reflect domain-specific judgment complexity. This work establishes a foundation for more trustworthy and accountable RAG systems in enterprise information-seeking applications, where detecting and resolving contradictions is essential for reducing risk and ensuring compliance.
I Introduction
Large language models (LLMs) are increasingly used across enterprises to draft policies, summarize compliance reports, generate HR documents, and support decision-making. These tasks differ from open-domain generation or short-form question answering. Enterprise text is dense, context-dependent, and often contains contradictions [^1] Contradictions may correspond to conflicting business rules, overlapping compliance guidelines, or inconsistent reporting., either introduced inadvertently or arising from evolving regulations, competing departmental priorities, or changes in strategy. In high-stakes enterprise settings, such as compliance reporting, financial filings, and policy drafting, these contradictions are more than nuisances. They can lead to regulatory violations, operational inefficiencies, unclear ownership of responsibilities, or costly business errors.
LLMs still struggle to reason over such conflicts or identify them reliably. Retrieval-augmented generation (RAG) pipelines aim to ground outputs in external evidence, but unresolved conflicts in retrieved documents like outdated policies and contradictory department guidelines often slip through unchecked. When LLMs answer enterprise queries without retrieval grounding, they hallucinate heavily, sometimes fabricating procedures or misrepresenting standards [1]. Even enterprise-domain RAG systems show high hallucination rates, with over 17% of benchmark queries producing flawed outputs due to irrelevant or conflicting evidence [2]. When contradictions in the input are not resolved, generation models tend to merge them, producing misleading or risky content.
To benchmark and mitigate these failures, we need controlled generation of contradictions. Unlike natural contradictions that occur by accident, we focus on intentional contradiction injection, i.e., generating enterprise-style documents that contain subtle, structured contradictions by design. These contradictions may occur within a document (self-contradictions) or across documents (pairwise contradictions). This controllability is essential for evaluating whether models can detect, resolve, or reason over contradictions when needed. It also enables robust stress-testing of RAG pipelines and contradiction-aware generation systems.
Contradiction detection remains difficult for both humans and machines. Many conflicts are subtle, context-dependent, or buried in long documents, and there is a scarcity of domain-specific datasets. Even state-of-the-art models like GPT-4 [3] and LLaMAv2 [4] perform only slightly better than chance on contradiction detection tasks [5]. These challenges are compounded by the difficulty of defining ground truth in enterprise corpora. Our framework incorporates a human-in-the-loop setup for quality assurance and for annotating contradiction pairs to support robust evaluation.
I-A Challenges in Contradiction Benchmarking
We highlight four major challenges that make contradiction analysis retrieval-augmented generation (RAG) both necessary and difficult:
  1. 1.
    Complex Enterprise Language: Existing contradiction benchmarks often use simplified sentences, whereas real enterprise documents (policies, compliance reports, HR manuals) contain jargon-heavy, multi-clause structures.
  2. 2.
    Limited Contradiction Types: Most datasets only handle binary contradiction/entailment, overlooking enterprise-specific cases like temporal mismatches, numerical inconsistencies, conflicting responsibilities, or strategy reversals.
  3. 3.
    Lack of Cross-Document Contradictions: Prior work emphasizes single-document conflicts, while in enterprises, contradictions frequently span multiple policies, reports, or department outputs.
  4. 4.
    Manual Contradiction Detection is Time-Intensive: Identifying contradictions across large corpora of policies, contracts, or filings requires tedious line-by-line checks, making manual detection impractical at scale.
I-B Our Contribution
To address these challenges, we introduce a novel multi-agent contradiction-aware benchmark generation framework that supports generation of realistic enterprise-style documents:
  1. 1.
    Realistic Enterprise Language: We generate synthetic documents that mirror enterprise tone, structure, and metadata. The framework supports customizable domains (e.g., HR, compliance, finance, strategy) and ensures quality through perplexity checks, LLM scoring, and human validation.
  2. 2.
    Rich Contradiction Taxonomy: We explicitly model six types of contradictions (Table I) going beyond simple negation to capture conflicts common in enterprise workflows, including temporal, numerical, and authority-based inconsistencies.
  3. 3.
    Pairwise Contradictions: We enable controlled creation of both self-contradictions (within a single document) and pairwise contradictions (across documents), simulating real-world enterprise conflicts between departments or policies.
  4. 4.
    Contradiction-aware Retrieval Evaluation: We provide an automated mining mechanism that detects contradictions using NLI models and LLM-based reasoning with confidence-weighted hybrid scoring. This is followed by human-in-the-loop validation through annotations conducted on the generated contradiction pairs.
  5. 5.
    Human-in-the-Loop Supervision: Human oversight is embedded at multiple stages, from verifying fluency and realism to resolving ambiguous contradiction cases, ensuring that generated corpora reflect enterprise-level fidelity and human judgment complexity.
To our knowledge, this is one of the first multi-agent frameworks for generating enterprise-style documents with controllable contradictions. The resulting synthetic corpora allow researchers to evaluate how well systems detect and reason over conflicts, providing a foundation for more reliable enterprise document generation, retrieval filtering, and contradiction-aware evaluation.
Contradiction Type Description Example
Temporal Contradicts date or time of an event. “Starts Jan 15” vs. “Starts end of Q1”
Numerical Conflicting numbers, values, or percentages. “\$12M surplus” vs. “\$5M deficit”
Authority Different source or issuer of a statement. “Issued by Compliance Office” vs. “Issued by Strategy Unit”
Process Conflicting procedures or operational routes. “Submit via HR portal” vs. “Submit through admins”
Policy Reversal One statement negates the other directly. “Remote work mandatory” vs. “Remote work not permitted”
Specificity One statement is more general or narrow than the other. “Applies globally” vs. “Applies only to APAC”
TABLE I: Taxonomy of Contradiction Types
II Background and Related Work
II-A Contradictions in Generation and Retrieval
Natural Language Inference (NLI) classifies the logical relationship between two sentences as entailment, neutrality, or contradiction. Contradiction detection focuses on identifying when two statements cannot both be true under the same interpretation. Contradictions appear either within a single document (self-contradictions) or across multiple documents covering related content (pairwise contradictions). In retrieval-augmented generation (RAG), unresolved conflicts like outdated policies, overlapping compliance guidelines, or inconsistent departmental reports can cause models to produce factually incorrect or logically inconsistent outputs. Left unfiltered, these contradictions represent a high-risk form of hallucination in enterprise settings. Consider the following example, where two HR policies specify different waiting periods for remote work eligibility.
Example: Contradictory Policies Input Documents: Policy A: “Employees are eligible for remote work after 30 days of service.” Policy B: “Employees are eligible for remote work after 90 days of service.” Query: “When can new employees start working remotely?” LLM Output: “New employees are eligible for remote work after 30 to 90 days, depending on circumstances.”
The model merges contradictory rules instead of resolving them, producing an ambiguous and non-compliant answer. ContraGen addresses this by intentionally injecting such conflicts in synthetic enterprise documents, enabling systematic evaluation of whether models can detect, localize, and resolve them.
Contradictions in RAG are increasingly studied as both hallucinations and knowledge conflicts, tied to broader challenges of factuality and reliability in LLM outputs. Zhang and Zhang [6] review hallucination mitigation in retrieval-augmented LLMs, showing that contradictions often stem from retrieval failures (outdated or conflicting sources) and generation deficiencies (context conflict, misalignment), making them one of the most risk-prone forms of hallucination in RAG. Wang et al. [7] propose a framework for evaluating knowledge conflicts between parametric and retrieved information, finding that while LLMs can detect conflicts, they struggle to localize them or provide distinct answers. Together, these studies suggest that contradictions in enterprise and compliance documents are not incidental errors but structural bottlenecks in retrieval and generation, motivating benchmarks that intentionally embed self and pairwise contradictions.
Despite growing interest, existing contradiction benchmarks are poorly suited for stress-testing generative systems. Early work in NLI introduced datasets like SNLI [8], MNLI [9], and ANLI [10], but these focus on isolated sentence pairs and lack the linguistic and structural complexity of enterprise documents. Recent datasets such as ContraDoc [5] and ECON [11] move toward document- and retrieval-scale evaluation. ContraDoc showed that even GPT-4 struggles with subtle internal inconsistencies in long documents, and trained annotators also miss many contradictions, highlighting that even human review is unreliable at scale. ECON distinguishes between factoid and answer conflicts, showing that models often default to internal priors when evidence is inconsistent. Gokul et al. [12] extend this line of work to synthetic RAG pipelines by generating controlled contradictions and testing LLMs as context validators, but they consider only three types: self, pairwise, and conditional contradictions.
These works highlight progress toward contradiction-aware evaluation but leave open the question of how to build realistic, domain-specific corpora that reflect the retrieval conflicts faced in enterprise applications—where overlapping reports, shifting KPIs, and conflicting compliance requirements are the norm.
Figure 1: ContraGen - Overall Workflow
II-B Multi-Agent and Generative Systems
Recent work highlights the growing role of multi-agent architectures in improving reliability and structure in LLM-based generation. Musumeci et al. [13] show that agent specialization—splitting tasks such as template decomposition, retrieval, and iterative drafting—can reduce hallucinations and produce more consistent public administration documents. Xie et al. [14] survey multimodal agents, emphasizing the need for perception, planning, and memory as core components of scalable deployment. Wang et al. [15] demonstrate that structured communication and hierarchical refinement in multi-agent systems can improve task performance in collaborative settings. M-RAG [16] leverages multi-agent reinforcement learning over database partitions, showing how specialization in retrieval and refinement can reduce noise and improve reliability.
Building on these insights, we design a multi-agent framework that applies specialization to contradiction-aware corpus generation. By decoupling generation, contradiction injection, and evaluation into separate roles, our approach provides finer control over the types of contradictions introduced and automates quality checks, directly addressing the gap in current benchmarks, which lack realistic, controllable corpora that reflect the retrieval conflicts common in enterprise and compliance workflows.
III System Architecture
ContraGen is a modular multi-agent system designed for generating and evaluating enterprise documents enriched with controlled contradictions. It supports document creation, contradiction injection, fluency validation, hybrid contradiction mining, and retrieval-based fact-checking. All agents communicate via Pyro4-based remote procedure calls and are coordinated by a central orchestrator. The overall architecture is shown in Figure 1.
ContraGen comprises three specialized agents, each addressing a distinct stage in the document generation and evaluation pipeline. The Content Generator Agent produces structured metadata and base documents for selected subdomains, injects self- and pairwise contradictions via instruction-tuned prompts, and validates fluency using relative perplexity scores while logging detailed metadata, including conflict pairs. The Contradiction Mining Agent employs a hybrid approach that combines a pretrained Natural Language Inference (NLI) model with an LLM judge, leveraging semantic similarity filtering to identify relevant sentence pairs and detect both self and cross-document contradictions. A human-in-the-loop complements this process by verifying that injected contradictions actually manifest as contradictions in context, manually adding injected cases that the system fails to detect, and annotating detections. Since injected conflicts can interact with surrounding content and create additional, unintended inconsistencies, human reviewers also provide contextual judgments to ensure that the evaluation set faithfully represents both intended and emergent contradictions. Finally, the Retrieval Verifiability Agent evaluates whether conflicting statements can be externally verified using publicly available evidence, enabling a two-track evaluation framework that distinguishes between retrieval-verifiable and retrieval-resistant contradictions.
Figure 2: Contradiction Generation
III-A Contradiction-Aware Content Generation Agent
III-A1 Metadata-Driven Base Document Generation
To initiate generation, we construct a realistic organizational profile that serves context for generating metadata. The scope of enterprise domain is organized into six categories. Given the organization description and subdomain structure, metadata is generated containing title, topic, date, department, location, document type, and authority level. This structured metadata serves as the source for multi-paragraph enterprise documents generated in assertive, policy-oriented language (see Prompt: Base Document Generation in Appendix A-A). We discuss the exact organization profile and domain-subdomain structure using in our framework in section IV.
III-A2 Perplexity-Based Fluency Control
Perplexity is a fundamental metric in Natural Language Processing used to assess the quality of language models and, by extension, to gauge the fluency and coherence of generated text. A language model that can accurately predict the next words in a sequence is more likely to generate fluent and grammatically correct text. Given a document represented as a sequence of tokens $w_{1},w_{2},\ldots,w_{N}$, the perplexity is:
$PP(w_{1:N})=\exp\left(-\frac{1}{N}\sum_{i=1}^{N}\log P(w_{i}\mid w_{<i})\right).$
where $P(w_{i}\mid w_{<i})$ is the probability assigned to the $i$-th token given its preceding context $w_{<i}$, and $N$ is the total number of tokens in the document. A lower perplexity indicates that the model finds the text more predictable (i.e., fluent) under its learned distribution. To ensure fluency preservation after contradiction injection, the perplexity of both the base and contradicted document is computed using pretrained language model GPT2 [17]. Let $PPL_{\text{base}}$ denote the perplexity of the base document and $PPL_{\text{contr}}$ the perplexity after contradiction injection. If the perplexity $PPL>22$, the generated document is rejected and re-tested. The relative change is defined as:
$\Delta_{\text{rel}}=\frac{PPL_{\text{contr}}-PPL_{\text{base}}}{PPL_{\text{base}}}$
III-A3 Contradiction Injection
To embed a self-contradiction (see Figure 2) , a target sentence is first selected using an LLM prompt that identifies a factual, specific, and authoritative statement. A second prompt with few-shot contradiction type examples generates a direct and confident contradiction statement, explicitly avoiding hedging language (e.g., “however,” “may,” “sometimes”). Finally, a third prompt blends the contradiction seamlessly into the document, preserving coherence, tone, and style. The prompts used are mentioned in Appendix A-B.
Pairwise contradictions are created by selecting a target statement from a source document $d_{1}$ and generating a corresponding contradictory statement as described earlier. The contradiction is embedded into a separate document $d_{2}$, ensuring stylistic and structural consistency while omitting the original target statement. This results in a document pair $\langle d_{1},d_{2}\rangle$, where $d_{2}$ contradicts a key assertion in $d_{1}$. The defined thresholds are validated, and any violation triggers document regeneration. This ensures fluent integration of contradictions while preventing unnatural degradation from language model nondeterminism.
$\Delta_{\text{rel}}^{\text{self}}\leq 0.05\quad\text{(self-contradiction, $\leq$ 5.5\% increase)}$ (1)
$\Delta_{\text{rel}}^{\text{pair}}\leq 0.075\quad\text{(pair contradiction, $\leq$ 7.5\% increase)}$ (2)
$PPL_{\text{contr}}\leq 22.0\quad\text{(absolute cap for sanity)}$ (3)
We conduct a human evaluation of generated documents containing injected contradictions using a 5-point Likert scale. The results show strong average ratings for fluency (4.2), specificity (4.7), and coherence (4.5), indicating the documents were well-written, detailed, and structurally consistent. Legitimacy scored slightly lower (3.5), suggesting that some documents lacked the authority expected of authentic enterprise-style texts. Contradictions were detected in only 54% of the cases, and when identified, were rated as subtle and realistic. Although our injected contradictions are naturalistic, they remain challenging for even human reviewers to reliably detect.
III-B Contradiction Mining Agent
Long-form documents may contain subtle inconsistencies which makes contradiction identification challenging for both humans and models. The Contradiction Mining Agent identifies high-likelihood contradiction pairs, which are then validated by human annotators. This agent ensures: 1) benchmark quality by verifying non-trivial, detectable contradictions and 2) provides a baseline for assessing RAG systems’ ability to recognize conflicts before generation.
NLI models are fast and widely used but they often struggle with spurious correlations, brittleness, and the difficulty of capturing nuanced reasoning. LLMs, on the other hand, offer richer reasoning but are computationally expensive. To identify contradictions at scale across documents, we combine fast retrieval-style filtering with multi-stage entailment reasoning. Our approach avoids exhaustive $\mathcal{O}(n^{2})$ sentence pair comparison.
  1. 1.
    Top-$k$ Semantic Filtering: For each source sentence, cosine similarity is computed against candidate targets using msmarco-distilbert-base-v3 [18], embeddings, fine-tuned for passage retrieval. The top-$k$ (k=5) pairs above threshold $\theta_{s}$ are retained, while short sentences (<10 words), numbers, and bullet points are filtered out. Duplicate pairs are removed via text-based hashing, effectively reducing the search space while preserving high-relevance candidates.
  2. 2.
    NLI-Based Contradiction Classification: The filtered sentence pairs are evaluated using a pretrained Natural Language Inference (NLI) model (facebook/bart-large-mnli) [19] to assign a label $\ell_{\text{NLI}}\in{\text{contradiction},\text{neutral},\text{entailment}}$ with confidence $p_{\text{NLI}}\in[0,1]$. Pairs labeled as contradictions or with $p_{\text{NLI}}\leq\theta_{\text{conf}}$ ($\theta_{\text{conf}}=0.7$) are forwarded for further assessment by a large language model (LLM).
  3. 3.
    LLM-Based Contradiction Judgment: For the retained pairs, GPT-4o [3] is prompted (see appendix A-C) to assess whether both statements can be simultaneously true. It returns a binary contradiction label $\ell_{\text{LLM}}\in{0,1}$, a justification, and a confidence score $p_{\text{LLM}}\in[0,1]$ reflecting its certainty.
  4. 4.
    Confidence-Weighted Hybrid Scoring: A hybrid contradiction score is computed as:
    $s_{\text{hybrid}}=w_{\text{NLI}}\cdot\ell_{\text{NLI}}+w_{\text{LLM}}\cdot\ell_{\text{LLM}},$
    where $\ell_{\text{NLI}},\ell_{\text{LLM}}\in\{0,1\}$ represent binary contradiction labels (1 = contradiction), and weights $w_{\text{NLI}},w_{\text{LLM}}$ are derived from the model confidences:
    $w_{\text{NLI}}=\frac{p_{\text{NLI}}}{p_{\text{NLI}}+p_{\text{LLM}}},\quad w_{\text{LLM}}=\frac{p_{\text{LLM}}}{p_{\text{NLI}}+p_{\text{LLM}}}$
    ensuring $w_{\text{NLI}}+w_{\text{LLM}}=1$. A pair is classified as contradictory if $s_{\text{hybrid}}>\tau$, where $\tau$ is a decision threshold ($\tau=0.5$).
Human-in-the-Loop (HITL) annotations are essential for ensuring both the realism of contradiction generation and the reliability of contradiction detection in long-form enterprise texts. Automated methods can inject and flag potential inconsistencies, but human expertise is required to validate plausibility and establish high-quality ground truth. For example, authority contradictions (e.g., Board approval vs. unilateral CEO approval) and cross-document conflicts (e.g., NDA permitting affiliate disclosure vs. master agreement prohibiting third-party disclosure) require expert interpretation to assess incompatibility. To quantify annotation reliability, we measured inter-annotator agreement (IAA) across both self-contradictions and pairwise contradictions. For self-contradictions ($n=141$), annotators achieved $95.0\%$ agreement with Cohen’s $\kappa=0.8973$ and Krippendorff’s $\alpha=0.8976$. For pairwise contradictions ($n=426$), the agreement was $96.7\%$ with Cohen’s $\kappa=0.8745$ and Krippendorff’s $\alpha=0.8746$. In practice, samples falling below a $90\%$ agreement threshold are adjudicated by a subject matter expert who selects the ground-truth label. Embedding such human judgment improves dataset fidelity, enhances model robustness, and ensures that contradiction detection remains interpretable and legally sound.
This staged framework enables interpretable contradiction detection in semi-structured documents, supporting both self- and pairwise analyses. Serving as a mining assistant rather than an oracle, it filters candidate pairs to a high-confidence subset for human verification, ensuring contradictions are both machine-detectable and human-confirmable while balancing scale and quality.
III-C Retrieval Verifiability Agent
A key challenge in evaluating contradiction resolution is distinguishing failure sources: retrieval gaps (missing evidence) versus reasoning gaps (requiring inference beyond retrieval). Existing benchmarks blur these distinctions, identification of RAG pipeline breakdowns. This agent addresses this gap by determining each contradiction pair $(s_{1},s_{2})$ is retrieval-verifiable or retrieval-resistant:
  • Retrieval-verifiable: Contradictions that can be resolved using reliable, publicly available evidence that (should be) retrievable by a RAG system.
  • Retrieval-resistant: Contradictions lacking sufficient supporting evidence, requiring reasoning or human judgment beyond retrieval.
This distinction makes contradiction evaluation diagnostic rather than merely descriptive. For example, consider two policy documents: “Employees are eligible for remote work after 30 days” vs. “Employees are eligible for remote work after 90 days.” If the official HR handbook specifies 90 days, the contradiction is retrieval-verifiable, indicating a retrieval weakness. If no authoritative source exists, it is retrieval-resistant, reflecting limitations in retrieval and the need for deeper reasoning. Labeling contradictions in this way localizes errors, making evaluation actionable for improving enterprise RAG systems.
IV Dataset Construction and Characteristics
To evaluate contradiction detection in a controlled yet realistic setting, a golden standard of 84 synthetic enterprise documents is constructed around a fictional multinational aerospace company which operates globally under strict confidentiality agreements: “a large aerospace company designing commercial aircraft and defense systems, with major manufacturing sites in Wichita, Toulouse, and Nagoya, and operating under strict international export controls.”
The corpus spans six high-level domains that reflect common organizational functions, each further divided into five subdomains:
  • HR and Personnel Management: HR Policy, Employee Relations, Talent Acquisition, Performance Management, Training and Development
  • Engineering and Development: Product Design, Software Development, Quality Assurance, R&D, Prototyping
  • Sales and Marketing: Lead Generation, Market Research, Campaign Management, Brand Strategy, Sales Enablement
  • Finance and Operations: Budgeting, Financial Reporting, Supply Chain, Procurement, Inventory Management
  • IT and Technical Support: Network Administration, Help Desk, Cybersecurity, System Integration, Cloud Services
  • Legal and Compliance: Contract Management, Regulatory Compliance, Litigation Support, Policy Drafting, Risk Assessment
After the base documents are generated using the above metadata and passing fluency checks, contradictions are systematically injected in a domain-specific manner. Each document in HR and Personnel Management, IT and Technical Support, and Legal and Compliance contains one self-contradiction, while an interleaving strategy within Engineering and Development, Finance and Operations, and Sales and Marketing modifies every second document to contradict the preceding one. This setup creates both intra- and inter-document conflicts, simulating realistic organizational scenarios. As shown in Table II, temporal and specificity dominate across contradiction types, with temporal inconsistencies prevalent in self-contradictions and specificity conflicts more common in pairwise cases. A single contradiction was counted under multiple categories if applicable as it can simultaneously be.
Although contradictions were injected by category, unintended conflicts (e.g., temporal inconsistencies) occasionally appeared. Human annotators tracked whether the injected contradictions were truly preserved as contradictions, or whether they blended smoothly into the text and no longer stood out. They also recorded whether these cases created detectable conflicts and whether the system correctly identified them, enabling a precise evaluation of robustness and sensitivity to subtle or spurious contradictions.
Contradiction Type Self Pairwise
Temporal 25 25
Numerical 7 6
Specificity 16 24
Policy Reversal 26 14
Authority 4 9
Process 10 11
Total 88 89
TABLE II: Distribution of contradiction types for self and pairwise contradictions.
V Results and Discussion
To evaluate contradiction detection in enterprise-scale documents, we implemented three complementary strategies: a Natural Language Inference (NLI) baseline, an LLM-based Judge, and a Hybrid model that combines the two. The NLI method offers scalability and precision but struggles to capture broader document context. In contrast, the LLM-based Judge provides richer semantic reasoning, yet remains vulnerable to prompt sensitivity, hallucination, and overconfidence in its outputs. The Hybrid approach balances these trade-offs by integrating NLI predictions with LLM judgments under a confidence-weighted scheme, yielding more stable and interpretable labels. Empirical results show that each detector identifies a partially overlapping subset of contradictions, underscoring the value of the Hybrid approach in achieving both breadth and reliability in contradiction mining. Since no single approach ensures an exhaustive coverage, a unified evaluation dataset was constructed:
  • Each candidate pair was uniquely defined by (doc1_chunk, doc2_chunk), normalizing whitespace and formatting for consistency. The union of all detector outputs ensured inclusion of every potential contradiction.
  • Ground-truth injected contradictions were manually added to the union to assess detector recovery of known cases.
  • Human annotators reviewed surrounding context and assigned binary labels (human_label) indicating the presence or absence of contradiction.
This process resolved inconsistencies, consolidated duplicates, and produced a unified gold-standard dataset comprising 141 self-contradiction pairs and 426 pairwise-contradiction pairs, with representative examples shown in Tables III and IV.
Target Statement Contradiction Paragraph LLM Pred Human Label
A critical change is the implementation of a centralized digital recruitment platform across all sites. The implementation of the decentralized recruitment system at our Wichita, Toulouse, and Nagoya facilities will continue. FALSE TRUE
Following the initial assessment, documentation and evidence compilation is mandated to be completed within 30 calendar days. Documentation and evidence compilation must be completed within 90 calendar days. TRUE TRUE
Our compliance with international export controls has occasionally delayed shipments. Despite stringent export controls, these have been flawlessly integrated without costs. TRUE FALSE
A key finding is that components produced in Nagoya offer up to a 15% cost reduction compared to Wichita. As of June 18, 2024, our analysis indicates significant cost variations between Wichita and Nagoya. FALSE FALSE
TABLE III: Examples of self-contradictions with LLM predictions vs. human labels.
Doc1 Chunk Doc2 Chunk LLM Pred Human Label
As of Feb 10, 2024, the wing design software implementation has been completed. The implementation of the wing design software has been delayed until March 20, 2024. TRUE TRUE
Conceptual Design Phase is initiated by senior aerodynamics engineer, Akira Tanaka. Conceptual Design Phase will involve avionics engineers, overseen by Dr. Aiko Takahashi. TRUE FALSE
All new aircraft models must undergo a five-phase quality assurance process. Prototyping for Next-Gen Avionics will be structured in four phases. FALSE TRUE
Quarterly audits will be conducted by Senior Compliance Analyst, Emma Dubois. Quarterly audits have been initiated by CFO Maria Fernandez. FALSE FALSE
TABLE IV: Examples of pairwise contradictions with LLM predictions vs. human labels.
Method Accuracy Precision Recall F1
NLI 41.1% 40.0% 76.7% 52.6%
Hybrid 90.1% 92.6% 83.3% 87.7%
LLM Judge 82.3% 79.7% 78.3% 79.0%
TABLE V: Performance on self-contradiction detection.
Method Accuracy Precision Recall F1
NLI 29.3% 12.9% 63.1% 21.4%
Hybrid 87.8% 57.8% 73.8% 64.9%
LLM Judge 72.1% 34.1% 89.2% 49.4%
TABLE VI: Performance on pairwise contradiction detection.
The quantitative evaluation through precision, recall, and F1 metrics reveals consistent trends in detector performance:
1. Self-contradictions are easier to capture. As outlined in Table V both the hybrid and LLM judge methods achieve high accuracies ($\sim$82–90%) and F1-scores ($\sim$79–88%). In contrast, the pure NLI detector struggles, highlighting its limitations in handling nuanced business-process contradictions.
2. Pairwise contradictions are substantially more challenging. Table VI shows that NLI collapses to near-random performance (accuracy 41%, F1 53%), underscoring its inability to generalize across documents. The hybrid detector outperforms the LLM judge (F1 64.9% vs. 49.4%), driven by its higher precision (57.8%), though this comes at the cost of recall (73.8%). It provides a balanced trade-off, while the LLM judge catches most true contradictions but also flags many non-contradictions.
3. Distinct precision–recall trade-offs. The hybrid consistently favors precision and stability, while the LLM judge favors recall, surfacing a wider range of subtle cross-document conflicts albeit with more false positives. This divergence is particularly useful diagnostically: hybrid filters provide a reliable baseline, while LLM reasoning exposes contradictions that might otherwise be overlooked.
The hybrid method achieves the best overall balance, reaching 87.8% accuracy and an F1 score of 64.9%, showing robustness in cross-document settings where subtle contextual shifts and entity references complicate detection. These results underscore that self-contradictions are easier to capture, whereas pairwise contradictions remain significantly harder, even for the hybrid model.
The strong performance gap between self and pairwise contradictions highlights an open challenge: while single-document consistency checks can be effectively addressed with current methods, multi-document contradiction detection requires more advanced techniques. Future work should explore approaches such as joint reasoning across documents or retrieval-aware prompting to bridge this gap. Strengthening this capability is particularly critical for enterprise applications like compliance monitoring, regulatory audits, and contract review, where surfacing subtle conflicts early can substantially reduce legal and operational risks.
VI Conclusion and Future Work
We present ContraGen, a framework for systematic contradiction generation and detection in long-form enterprise documents, moving beyond traditional sentence-pair benchmarks. The framework creates, annotates, and evaluates contradictions at the document level, enabling rigorous assessment of both intra-document and cross-document consistency. This design addresses a critical gap in enterprise AI, where contradictions frequently arise across business processes, policies, and regulatory documents. Empirical results indicate that LLM as a judge provides the most balanced trade-off between precision and recall for both self-and pairwise-contradictions, yet detecting cross-document contradictions remains a significant challenge due to contextual variability across heterogeneous sources.
Limitations. This work faces several limitations. (1) computational and API expenses scale directly with the volume of generated documents and injected contradictions, making large-scale experimentation costly; (2) the evaluation relies on synthetic yet realistic enterprise-style documents, since confidentiality and compliance constraints restrict access to proprietary corpora and large-scale public datasets are not yet available. (3) establishing reliable ground-truth labels continues to depend on human annotators, limiting full automation; (4) although the framework applies thresholds (e.g., perplexity caps, similarity filters, decision cutoffs), we did not systematically tune parameters, limiting reproducibility across computational settings. Finally, this study represents an initial step toward controllable contradiction generation and detection, laying the groundwork for future comprehensive LLM benchmarks.
Future Work. Future research will pursue several directions: (1) developing benchmark suites that integrate synthetic and semi-real documents (e.g., public policy manuals, contracts) to broaden generalizability; (2) enhancing cross-document contradiction detection through retrieval-aware prompting, multi-hop reasoning, and systematic evidence alignment across heterogeneous sources such as contracts, compliance reports, and regulatory filings; (3) conducting rigorous ablations and sensitivity analyses, including parameter tuning for similarity thresholds, NLI backbones, and LLM judges, to improve reproducibility; (4) evaluating the robustness of RAG pipelines when confronted with conflicting enterprise evidence, with attention to whether failures stem from retrieval gaps or reasoning gaps; and (5) considering practical guardrails for moving from synthetic to real-world corpora, including redaction, access control, and auditing mechanisms to ensure privacy and compliance. Together, these steps will help scale ContraGen from a controlled benchmark to a deployable evaluation framework for enterprise RAG systems.
Appendix A Appendix
A-A Prompts for Data Generation
Prompt: Base Document Generation Generate a professional, coherent business document (4–6 paragraphs) for {company} in the domain of {domain}. The topic is ‘{topic}’, related to ‘{phrase}’. Use the format of a {doc_type}. Structure the document with 4–6 clearly titled sections, paragraphs, numbered lists, or bullets as needed with realistic internal formatting. Include at least one CLEAR, FACTUAL, ASSERTIVE sentence early in the document - ideally in paragraph 1 or 2. This sentence will be contradicted later, so make it specific and important. Do not add contradictory information yet. The content should create dates and time intervals close to {date} and associate them with any requirements. The content should create specific personnel titles and contact info for {department}. The content should create specific sub-locations within {locations}. Vary the audience tone between internal staff, legal reviewers, or external partners. Use diverse names, avoiding placeholders like “John Smith”. Include names appropriate to a global workforce. Be assertive: state policies directly. Avoid vague phrases like “such as” or “for example”. If a process is mentioned, describe it in detail (e.g., steps, links, roles, submission method). At the end, include: NEW PEOPLE META DATA: Names, roles, and affiliations of people mentioned in the document. NEW DOCUMENT META DATA: Titles and dates of referenced documents (e.g., forms, policy memos, review notices).
A-B Prompts for Contradiction Injection
Prompt: Identify Statement You are analyzing a business document to identify the most important/specific statement for contradiction injection. From the following document, select ONE sentence that is: Clear, factual, and assertive Important to the document’s main topic Specific (contains numbers, dates, procedures, or concrete details) Return ONLY the selected sentence, nothing else. DOCUMENT: {document_text}
Prompt: Generate Contradiction You are helping generate business documents that contain realistic internal contradictions. You will be given the original document and one important sentence selected from it (the target statement). Your task: Write 1 or 2 sentences that DIRECTLY CONTRADICTS the target statement after selecting one of the contradiction types mentioned in the few-shot examples. The contradiction should be HARD and DIRECT, not soft or hedging. Use the same tone and style as the original document. Do NOT use words like ‘however’, ‘while’, ‘although’, ‘but’, ‘may’, ‘might’, ‘could’, ‘sometimes’, ‘certain’, ‘extended’, ‘flexibility’. Make a clear, confident business statement that contradicts the target. The contradiction should be realistic, specific, and plausible within a formal enterprise setting. The paragraph should sound like a natural part of the document, not an exception or explanation. FEW-SHOT EXAMPLES: TARGET: The new performance evaluation cycle begins on January 15, 2024.
CONTRADICTION: The official launch of the new evaluation cycle will be in late Q1 accommodating inter-departmental calibration sessions.
TYPE: temporal
TARGET: The North America team exceeded Q2 revenue targets by \$12M.
CONTRADICTION: The North America team fell short of Q2 revenue by \$5M.
TYPE: numerical
TARGET: This policy is issued under the authority of the Global Compliance Office.
CONTRADICTION: This policy is issued by the Strategy Operations Unit.
TYPE: authority
TARGET: All employee leave requests must be submitted via the HR portal.
CONTRADICTION: All employee leave requests must be submitted through departmental admins.
TYPE: process
TARGET: Remote work is mandatory for Tier 1 engineers.
CONTRADICTION: Remote work is not permitted for any Tier 1 engineering staff.
TYPE: policy reversal
TARGET: The new training policy applies to all regional offices globally.
CONTRADICTION: The new training policy applies to the APAC offices.
TYPE: specificity
TARGET STATEMENT: {target_statement} ORIGINAL DOCUMENT: {document_text}
Prompt: Blend Contradiction into Document You are an expert business writer. You are given a business document and a contradiction paragraph. Your task is to rewrite the document, blending the contradiction paragraph into the most natural place in the document, ideally away from the original target statement. Instructions: Keep the original target statement in the document. You may rewrite or rephrase the document as needed, but the contradiction must be present and subtle. The final document should flow naturally, with the contradiction paragraph integrated realistically and not artificially or as an explanation. Use the same tone and style as the original document. The contradiction must be CLEAR and FACTUAL - do not explain, soften, justify, or hedge. Do NOT use words like ‘however’, ‘while’, ‘although’, ‘but’, ‘may’, ‘might’, ‘could’, ‘sometimes’, ‘certain’, ‘extended’, ‘flexibility’. Do NOT include explanations, reasoning, timeline adjustments, or supporting logic. Just state the conflicting fact. At the end, include: NEW PEOPLE META DATA: Names, roles, and affiliations of people mentioned in the document. NEW DOCUMENT META DATA: Titles and dates of referenced documents (e.g., forms, policy memos, review notices). DOCUMENT: {base_content} ORIGINAL TARGET STATEMENT: {target_statement} CONTRADICTION PARAGRAPH: {contradiction_paragraph} Return ONLY the final, blended document.
A-C Prompt used for LLM Contradiction Judge
Prompt: LLM Contradiction Judge You are an expert at detecting logical contradictions in text. Analyze the following two sentences and determine if they contradict each other. Sentence 1: “{sentence1}” Sentence 2: “{sentence2}” Consider: 1. Do these sentences make opposing claims about the same subject? 2. Are they logically incompatible? 3. Could both be true in the same context? 4. Are they discussing different aspects of the same topic? Respond with a JSON object containing: "contradiction": true/false "reasoning": brief explanation of your decision "confidence": float in [0.0–1.0] representing certainty Example response:
{"contradiction": false,
 "reasoning": "Both sentences discuss
 different aspects of compliance
 without conflicting",
 "confidence": 0.8}